### v2

# GPTKB: Building Very Large Knowledge Bases from Language Models

Abstract
General-domain knowledge bases (KB), in particular the “big three” –Wikidata, Yago and DBpedia– are the backbone of many intelligent applications. While these three have seen steady development, comprehensive KB construction at large has seen few fresh attempts.
In this work, we propose to build a large general-domain KB entirely from a large language model (LLM). We demonstrate the feasibility of large-scale KB construction from LLMs, while highlighting specific challenges arising around entity recognition, entity and property canonicalization, and taxonomy construction. As a prototype, we use GPT-4o-mini to construct GPTKB, which contains 105 million triples for more than 2.9 million entities, at a cost 100x less than previous KB construction projects.
Our work is a landmark for two fields: For NLP, for the first time, it provides constructive insights into the knowledge (or beliefs) of LLMs. For the Semantic Web, it shows novel ways forward for the long-standing challenge of general-domain KB construction. GPTKB is accessible at http://gptkb.org.
GPTKB: Building Very Large Knowledge Bases from Language Models
Yujia Hu1   Shrestha Ghosh2   Tuan-Phong Nguyen2   Simon Razniewski1 1ScaDS.AI & TU Dresden, Germany 2Max Planck Institute for Informatics, Saarbrücken, Germany yujia.hu@mailbox.tu-dresden.de  {ghoshs,tuanphong}@mpi-inf.mpg.de  simon.razniewski@tu-dresden.de
1 Introduction
Figure 1: Overview of our approach for LM-based KB construction.
General-world knowledge bases (KB) like Wikidata Vrandecic and Krötzsch (2014), Yago Suchanek et al. (2007) and DBpedia Auer et al. (2007) are important backbones for intelligent applications. While these projects exist for over a decade, innovation in the field of general-world KB construction is low, with neither fundamental paradigm shifts nor major novel projects emerging Weikum et al. (2021).
Recently, large language models (LLMs) stirred up many fields of AI Bubeck et al. (2023), and also been proposed as sources for structured knowledge Petroni et al. (2019). More specifically, Cohen et al. (2023) showed how, in principle, one can build a KB from an LLM by simple factual prompts and iterative graph expansion, but have not attempted this at scale.
The success of LLMs has also raised important intrinsic questions, in particular, what and how much these models know or believe Petroni et al. (2019); Jiang et al. (2020); Roberts et al. (2020); Veseli et al. (2023); Sun et al. (2024); Wu et al. (2024b). A large set of benchmarks and studies investigate this via example-based prompting, e.g., to determine, how many answers to common benchmarking question answering (QA) datasets are known by LLMs. However, all these works remain non-exhaustive, investigating samples from specific datasets or domains, without attempting to materialize all knowledge of an LLM.
In this paper we propose to comprehensively materialize the knowledge/beliefs [^1] The terminology here is contentious, see Section 6.1. of LLMs into a KB. In particular, we propose to use iterative graph expansion to obtain the complete set of (named-entity-centric) LLM knowledge, and to consolidate that knowledge with LLM-based entity disambiguation, class and relation canonicalization, and taxonomy induction (see Fig. 1). This proposal faces several challenges:
  1. 1.
    Termination/cost/runtime: State-of-the-art KBs contain millions of entries, and LLMs are known to hallucinate. It is therefore not clear if and when iterative graph exploration will terminate, and how to perform it under practical monetary and time constraints.
  2. 2.
    Flexible and relevant knowledge elicitation: LLMs possess a wide variety of knowledge, and the amount of knowlege per entity varies. We need a method that elicits as much general-world knowledge as possible, without encouraging hallucinations, and without falling into bottomless corners, e.g., open-ended phrases or translations.
  3. 3.
    Canonicalization and disambiguation: Hallmarks of existing KBs are disambiguated entities and coherent classes, relations and taxonomies. Iterative prompting risks surfacing expressions that are not globally coherent.
Our approach builds on the following ideas: To overcome scaling issues, and obtain relevant knowledge, we utilize a commercial API that allows to massively send batch requests (GPT-4o-mini), and utilize named entity recognition (NER) and carefully crafted prompts to restrict the explored spaces, along with prompts that allow varied answer sizes. To obtain a coherent KB, we perform a set of canonicalization and disambiguation steps, entirely relying on the LLM itself. In summary, our salient contributions are:
  1. 1.
    To the best of our knowledge, we are the first to propose to construct comprehensive KBs entirely from LLMs.
  2. 2.
    We develop a KB construction pipeline that overcomes termination, quality and runtime issues.
  3. 3.
    We construct GPTKB, the first large-scale KB entirely built from an LLM, containing over 104M assertions for over 2.9M entities.
Our work is a significant step for two communities: For the NLP community, for the first time, we provide a proof-of-concept methodology that enables constructive insights into what LLMs know (or believe). For the Semantic Web community, we provide fresh momentum for the long stale task of open-domain KB construction, and provide code, a concrete resource, GPTKB, both as a 3.8 GB download, and via an online browsing interface and SPARQL query interface at http://gptkb.org.
2 Related work
Large-scale KB construction
Dominating public large-scale KBs are Wikidata Vrandecic and Krötzsch (2014), Yago Suchanek et al. (2007) and DBpedia Auer et al. (2007), all started more than 10 years ago. While Wikidata is constructed by volunteers, Yago and DBpedia represent the paradigm of (semi-)structured information harvesting and integration, extracting in particular from Wikipedia infoboxes and Wikidata Weikum et al. (2021). They all remain incomplete Razniewski et al. (2024), warranting the search for novel paradigms. Commercial projects like the Google KG Singhal (2012) or Amazon’s KG Dong et al. (2020) have usually followed these approaches. By comparison, text-based KB construction, e.g., in NELL Mitchell et al. (2018) or ReVerb Fader et al. (2011), has achieved less adoption. Our approach is more related to the latter approaches, as LLMs are distillations of large text corpora. Table 1 gives an overview of major KB projects.
KB #entities #assertions
Wikidata 113M 1.62B
Wikidata5m 5M 20M
Yago 4.5 50M 140M
DBpedia 3.8M 75M
NELL ? 12M
ReVerb ? 15M
GPTKB 2.9M 105M
Table 1: Size comparison1 of major KBs. The ones in the top block are Wikimedia-related, NELL and ReVerb are text-extracted.
11footnotetext: Sources: Wikidata: See https://www.wikidata.org/wiki/Wikidata:Statistics and https://grafana.wikimedia.org/d/000000175/wikidata-datamodel-statements. Wikidata5m: Wang et al. (2021). Yago4.5: Suchanek et al. (2024). DBpedia: English version as per Table 2 in Lehmann et al. (2015). NELL: As per Figure 5 (left) in Mitchell et al. (2018). Reverb: As per Lin et al. (2012) and http://web.archive.org/web/20220307185343/https://openie.allenai.org/
Extent of LLM knowledge
Since the emergence of LLMs, the question of how much these models know has frequently been raised Roberts et al. (2020); Jiang et al. (2020); Veseli et al. (2023); Sun et al. (2024); Wu et al. (2024b). So far, the widely adopted approach is to sample a domain of interest, e.g., question-answer pairs from a common benchmark, or triples from Wikidata, probe the LLM, and compute the fraction that the LLM can correctly answer/complete. Yet this does not provide an overall estimate of LLM knowledge, as the LLM might hold a lot of knowledge outside the focus of existing benchmarks. For instance, we found that GPT-4o-mini holds substantial knowledge about people’s hobbies, which are not well covered in existing KBs.
Knowledge extraction from LLMs
Most previous works only transiently probe LLMs for individual objects of subject-relation-pairs like Vannevar_Bush-citizenship Petroni et al. (2019); Veseli et al. (2023); Sun et al. (2024); Wu et al. (2024b), and use samples to estimate the LLMs knowledge. Funk et al. (2023) explore ab-inito taxonomy construction from LLMs. Kassner et al. (2021) propose a small-scale persistent memory component for ensuring LLM answers remain consistent over multiple prompts. Nguyen et al. (2024) iteratively prompt GPT-3.5 for obtaining 167K sentences containing cultural commonsense for 11K subjects. Cohen et al. (2023) propose to iteratively prompt GPT-3 for relations and relational assertions for triples. In difference to Nguyen et al. (2024), we are after structured content (triples), at a much larger scale. In difference to Cohen et al. (2023), we unveil the practical challenges that such a graph exploration faces, tackle them, and perform LLM-based KB construction at scale.
3 Methodology
An overview of our approach is shown in Fig. 1. In the first phase, we iteratively elicit LLM triples for a given subject, and enqueue newly encountered named entities for further triple elicitation. In the second phase, we consolidate the resulting triple set by canonicalizing entities, relations and classes, and by constructing a coherent taxonomy. Our paradigm is that as much knowledge as possible should come from the LLM itself, we therefore refrain from imposing any standardized vocabularies, or using existing KBs as disambiguation references.
3.1 Knowledge elicitation
Iterative graph expansion
We start the extraction process from a single seed subject, Vannevar Bush, the visionary behind the concept of hyperlink-based knowledge organization Bush (1945). From the triples obtained for him, just like in web crawling, we can then identify further entities (e.g., MIT (affiliation) or Everett, MA (birth place)), for which we can then elicit further knowledge, and so on Cohen et al. (2023). Figure 2 illustrates how in 3 hops, we arrive at entities of diverse types, such as historical event (Boston Tea Party, Manhattan Project), newspaper (The Times) and magazine (The New Republic).
Figure 2: Graph exploration from seed entity Vannevar Bush.
Knowledge prompting
A major challenge in knowledge elicitation is to elicit as much knowledge as possible, but at the same time, not encouraging hallucinations. We found that without guidelines on the expected number of triples, LLMs systematically returned much fewer triples than they could correctly return. However, a static number would not be suitable either, as e.g., for Albert Einstein, an LLM should return many more triples than for the average person. We solve this via a subtask where the LLM first specifies the number of triples it believes to know for the subject, with two ranges as orientation. In difference to Cohen et al. (2023), we also streamlined the process by dropping the separate elicitation of relations and relation-specific triples. To structure the knowledge, we also request that at least one instanceOf triple should be returned. Output parsing is eased by using the structured outputs feature of OpenAI’s API. In turn, this also reduces hallucinations, e.g., additional qualifiers, textual descriptions, or similar. The full prompt is presented in Fig. 3.
You are a knowledge base construction expert. Given a subject entity, return all facts that you know for the subject as a list of subject, predicate, object triples. The number of facts may be very high, between 50 to 100 or more, for very popular subjects. For less popular subjects, the number of facts can be very low, like 5 or 10. Important: - If you don’t know the subject, return an empty list. - If the subject is not a named entity, return an empty list. - If the subject is a named entity, include at least one triple where predicate is “instanceOf”. - Do not get too wordy. - Separate several objects into multiple triples with one object.
Figure 3: Our prompt for knowledge elicitation.
Named-entity recognition (NER)
Our early attempts at graph exploration were plagued by topical runaway into linguistic knowledge, translations, etc. To keep the process on target, we use spaCy’s NER module to enqueue only objects who are among common classes of named entities. We include language-specific NER models for 10 major languages, and employ a lightweight rule-based filter to exclude websites URLs and other noise. The NER component appears to still hold room for improvement, likely due to the fact that NER models are traditionally trained on sentences, while in our task, the inputs are only the short, context-free entity labels.
3.2 Knowledge consolidation
Because input prompt size directly relates to the complexity and cost of prompting, it is not possible to feed a large amount of existing entity or taxonomy knowledge into the knowledge elicitation phase. Consequently, the output exhibits a huge degree of redundancy and variance, which we consolidate post-hoc.
Relation clustering
The naive process generates 688,855 distinct relation names, with many obvious duplicates, e.g., instanceOf, isA, or InstanceOf. To remove redundancies within this set, we apply greedy clustering, based on embedding similarity computed via SentenceTransformers Reimers and Gurevych (2019). Specifically, we compare each relation $R$ with all more frequent ones, starting from the most frequent one, and merge relation $R$ into relation $S$, if the textual embedding similarity between the names of $R$ and $S$ is greater than:
$1.2\times\frac{\log{(\textit{freq}(R))}}{\log{(\textit{freq(most\_freq\_relation)})}}.$
In particular, the formula is designed so that the similarity threshold varies with the frequency of the relation, leading to a more aggressive removal of relations that occur with low frequency. This way, we reduce the number of relations to 2,133 (a 323x reduction).
Class clustering
The naive process also generates 83,770 distinct class names (objects for instanceOf), with both obvious duplicates, and many overly specific cases. We employ the same greedy clustering as for relations, with the same multiplier 1.2 in the merging criterion. This way, we reduce the number of classes to 367.
Taxonomy construction
The classes in the KB so far do not form a coherent, or even connected taxonomy, as the knowledge elicitation process only expands named entities. But we also cannot simply prompt for a complete taxonomy to put on top (e.g., following the approach in Funk et al. (2023)), because we need to include the existing classes. We chose to proceed in three steps: In step 1, we prompt the LLM to generate an initial high-level taxonomy of about 50 classes (see prompt in Fig. 4 (top)). In step 2, we prompt the LLM to assign to each class in the KB a score describing how general it is (see prompt in Fig. 4 (upper middle)), then use that score to rank the classes from most general to most specific. In step 3, we insert classes into the taxonomy, one at a time. We initially tried this with the full taxonomy as input, but found that the LLM struggled. We therefore now proceed recursively. Starting from the nodes below the root, we ask the LLM repeatedly, whether the given class to insert belongs below one of them, in which case we repeat this process recursively with that node’s children (see prompt in Fig. 4 (lower middle)). The recursion stops either when the given class is not a subclass of any child anymore (in which case it is inserted as sibling), or when a leaf is reached (in which case it is inserted as new child). Then we prompt LLM to update the (sub)taxonomy from current node with the given class (see prompt in Fig. 4 (bottom)). We feed the final taxonomy back into GPTKB via the rdfs:subClassOf relation. Further LLM-based refinement in the style of Peng et al. (2024) could be considered.
You are a knowledge base construction expert. Your task is to initialize a seed taxonomy with general categories, which you will update later with given classes. Please return only the seed taxonomy in json form with indentation: Class: <new class> You are a knowledge base construction expert. Your task is to create a taxonomy for a knowledge base. Beforehand, you need to give each given class a score describing how general it is. The score is a integer ranging only from 1, for the most general concept, to 10, for the most specific concept. Please return only the score of the given class: Candidate branches: <branches of current node> Class: <class to add> You are a knowledge base construction expert. Your task is to integrate a given class into the taxonomy. If the given class is a subclass of one of the candidate branches, return only the exact name of that branch. Otherwise, return only NULL. Taxonomy: <taxonomy from current node> Class: <class to add> You are a knowledge base construction expert. Your task is to update the given taxonomy with the given class. You can consider the categorization of the taxonomy, but you can not modify the names of the classes in the taxonomy. Please return only the updated taxonomy in json form:
Figure 4: Prompts for seed taxonomy construction (top), class generability scoring (upper middle), recursive class insertion check (lower middle), and (sub)taxonomy update (bottom).
Entity deduplication
Naive graph exploration frequently yields duplicates, for instance, John_F._Kennedy and John_Fitzgerald_Kennedy. To remove these duplicates without requiring extraordinary runtime, we follow the standard blocking-based deduplication approach Köpcke and Rahm (2010). This technique is usually based on choosing an appropriate blocking key: one attribute by whose values we obtain meaningful partitions on entities. The key should meet two criteria: on the one hand, it should be a necessary condition for equivalence; on the other hand, it should be distinctive enough to partition the entities into not too large blocks.
We focus on entities that are instances of the most interesting class, humans. We block entities by birth dates. [^2] Only considering those with a birth date at a resolution more specific than a year. Within each block, we consider an entity pair as duplicates if both their labels exhibit at least 0.85 similarity using SentenceTransformers, and 30% of their triples are exactly the same. We then merge these entities, keeping the name of the entity that was created first. In total, this step removes 6,987 duplicate instances of Person. More advanced methods could utilize LLMs themselves for deciding on entity equivalence Ding et al. (2024).
4 Implementation
LLM choice and parallelization
We chose GPT-4o-mini OpenAI (2024) for our experiments, because of its relatively good tradeoff between performance and cost, and its ability to process requests in batches. This means we can send multiple subjects from our queue at the same time, after startup, typically 5,000 entities per batch, and up to 100 batches in parallel. The model’s size is not publicly released, but has been estimated to be around 8B. [^3] See https://explodingtopics.com/blog/gpt-parameters, which in turn links to https://techcrunch.com/2024/07/18/openai-unveils-gpt-4o-mini-a-small-ai-model-powering-chatgpt, which references a statement by OpenAI that we could not find. But the underlying reasoning based on performance comparison with open models appears sensible.
Seed entity, result size, runtime and cost
We start the process from Vannevar Bush, the visionary of interlinked knowledge management. We require a total of 2,200 batches to prompt GPT for 5.8M entities, obtaining in the answers triples for over 2.9M entities of them, which took us about 27 hours. The subject queue was still not empty at that point, but contained overwhelmingly duplicate or hallucinated subjects (see discussion below). Including trial runs, constructing GPTKB costed us a total of \$2,200 for OpenAI API calls.
GPTKB statistics
Our KB contains a total of 105M triples for 2.9M entities, organized into 2,133 relations and 367 classes. This gives an average of 36 triples per subject, with two distinct clusters, in particular, 651K entities with 10 triples and 86K with 50 triples, with most others near these values. Of all triples, more than 37M have an entity as object, and more than 67M have a literal as an object. The average length of entity labels is 24.5 characters. Example output for Vannevar Bush is shown in Table 2.
Dataset provision and license
We provide our KB both as a download (3.8 GB in TTL format), via a web browsing interface, and via a SPARQL query endpoint at http://gptkb.org. We provide our KB under the noncommercial CC BY-NC 4.0 license, not out of commercial self-interest, but because we want to avoid potential misuses (see Section 8). This license is compatible with the permissive terms of use of OpenAI. [^4] https://openai.com/policies/sharing-publication-policy
5 Analysis
5.1 GPTKB content
The most frequent classes generated for entities are shown in Fig. 5, where we find that patent and Person dominate the dataset. Several more classes are actually subclasses of Person, e.g., Noble, Duke, historical figure. Including these, the total number of humans is close to 600K. We also find a largely hallucinated class, US Coast Guard cutters, among the top-3.
[…]
Figure 5: The 12 most frequent classes in GPTKB, which make up 32% of all entities.
Relation Object
instanceOf person
birthPlace Everett, Massachusetts
born March 11, 1890
affiliation Massachusetts Institute of Technology
United States government
influencedBy John von Neumann
knownFor Memex
awards National Medal of Science
Presidential Medal of Freedom
Table 2: Excerpt from 41 triples for subject Vannevar Bush in GPTKB. Entity objects for which further triples were found are highlighted in italics.
The most frequent properties in the whole KB are patentCitation (3.15M) and instanceOf (2.96M). with similarly generic properties following at the KB level. At the class level, we find more specific properties, for instance, the most frequent properties in the class Person are instanceOf (251K), hasOccupation (126K), knownFor (119K), nationality (114K).
We also analyze the geographic bias of our KB, using the listed nationalities as proxy. The results are shown in Fig. 6. We can observe a clear bias towards English-language nationalities, stronger than, e.g., in Wikidata Shaik et al. (2021), and a curious lack of others (e.g., Chinese only 2K), likely reflecting the (undisclosed) English-centric training corpus of GPT-4o-mini. Further interesting observations concern the LLM’s reluctance to express the gender property (only 30K triples, with female at 15K more present than male at 8K), likely reflecting the measures taken towards hiding sensitive attributes, and active debiasing.
[…]
Figure 6: Distribution of values for nationality.
5.2 Wikidata comparison
We compare with Wikidata on several aspects: First, we compute the fraction of subjects that exist in Wikidata. For a random sample of 200 subjects from GPTKB, 24% have an entity with exactly matching label in Wikidata. A further 6.5% have a non-empty search result, i.e., an entity of paraphrased or similar label. The remaining 69.5% appear novel. Second, we exemplarily look at the 41 triples for Vannevar Bush, of which we find that more than 10 are not contained in Wikidata, e.g., his affiliation with the US government, his children count (incorrect by 1), or him inventing the concept of hypertext. [^5] An exact comparison is not straightforward, because relation names do not perfectly align, and objects in GPTKB are substantially more wordy. Third, we identify several properties not modelled at all in Wikidata, for instance, historicalSignificance (270K triples), hobbies (30K triples), or hasArtStyle (11K triples). This indicates that GPTKB potentially contains a significant amount of novel knowledge. A more comprehensive KB comparison in the style of Färber et al. (2018) is planned.
5.3 Precision and hallucinated content
Evaluating the precision of large-scale KBs is not straightforward, because they contain a significant amount of long-tail knowledge, for which finding evidence or counter-evidence is difficult. In line with previous work Suchanek et al. (2007), we evaluate precision within the context of web-retrievable information.
In particular, for a given triple, we retrieve the top-5 web-page snippets for the search term <subject object> from a search engine API, then use another LLM, namely Llama-3.1 llama31, to judge whether the given triple appears true/plausible/implausible/false, given the snippets. We adapt the prompt used by Adam and Kliegr (2024), extending it with the “false” level. From a random sample of 200 triples, we find that, based on the snippets, 22.5% appear true, 57.5% appear plausible, 1% appear implausible, 19% appear false. We also sample 200 triples from entities that are instances of Person, finding there that 21.5% of their triples appear true, 52.5% appear plausible, 0% appear implausible, 26% appear false. Extrapolating the former results, we can estimate that GPTKB contains approximately 24M triples that can be web-verified as true, 60M that appear plausible according to web context, 1M that appear implausible, and 19M that are contradicted by web context.
In terms of subjects, we observed a wide variance. While for some classes, e.g., Person or Company, most entities appeared real, others showed a strong runaway, e.g., patents, where most codes were made up (e.g., US_04,500,002_A1), or United States Coast Guard Cutter (the LLM continuously added varying incorrect type numbers to real names, e.g., USCGC_Biscayne_(WLB-201)). We therefore also evaluate the web verifiability of subjects alone. Similar to the approach used for triples, we employ a search engine to retrieve 5 snippets, now given only the entity’s name as query. We then evaluate three levels: (i) The snippets contain an entity of exactly that name, (ii) the snippets make it likely that the entity exists, perhaps under a slightly different label, (iii) the snippets give no support for the existence of that entity. From a random sample of 200 entities, we find that 41.5% fall into the first category, 43.5% into the second, 15% into the third, implying an extrapolated 1.2M verifiable, 1.3M plausible, and 0.4M questionable KB subjects. Among subjects classified as Person, 67% fall into the first category, 21.5% into the second, and 11.5% into the third. In the contrast, only 0.5% of subjects under United States Coast Guard Cutter fall into the first category, 13.5% into the second and 86% into the third.
About the accuracy of the taxonomy, we follow the per-edge evaluation scheme proposed by Bordea et al. (2016), finding that 64% of all subclass-superclass edges are considered correct, using another LLM as judge. We also evaluate whether a superclass is most appropriate, by offering all siblings as alternatives. Here, for 70% of all subclass-superclass edges, the superclass is considered the best alternative. Note that the structure of our taxonomy still leaves room for improvement in terms of long-range dependencies and distributions, that are not easy to quantify or address locally.
6 Discussion
6.1 Lessons for LLM epistemiology
LLMs and the way they store, retrieve, and generate knowledge are under intense scrutiny from computational, as well as cognitive and philosophical perspectives Nosta (2024). We have shown that one can iteratively materialize internal knowledge of an LLM at an unprecedented scale, but the notion and definition of LLM knowledge itself is controversial Fierro et al. (2024). We adopt here the term knowledge base because it is common in the field, but would consider the term belief base equally appropriate. In terms of the definitions of knowledge that are introduced by Fierro et al. (2024), our output falls into the minimalistic sui-generis (g-knowledge) category, as our KB may contain false as well as inconsistent triples.
We can now, in principle, relate the number of parameters of an LLM to the amount of entity-centric triple knowledge it holds: As discussed in Section 4, GPT-4o-mini likely has in the order of 8B parameters. Given the 105M triples that we obtained, that makes 76 parameters per triple. [^6] From Allen-Zhu and Li (2024), a theoretical lower bound of 4 parameters/triple can be deduced, based on Remark 4.4, and their synthetic dataset containing 6 triples per entity. However, the generalizability of results on that synthetic dataset is unclear. We emphasize that this number captures only encyclopedic knowledge, yet that LLMs also possess other knowledge, e.g., linguistic knowledge, procedural knowledge, etc.
As already observed by Petroni et al. (2019), prompt formulations heavily influence LLM responses, and hence may bias the resulting KB. GPTKB presents one way of materializing LLM knowledge into structured format, but it is important to observe that LLM knowledge is inherently probabilistic, while our output (and most web-scale KBs) are binary. Any process translating from probabilistic to binary implicitly encodes a choice between precision and recall, and we do so by the triple counts given as guidance in the elicitation prompt. In our view, such a biasing is inherently unavoidable, and also not solved by instruction-free sample-only (few-shot) prompts as proposed by Wu et al. (2024b), as also examples bias the responses, both in terms of quantity, topics, and vocabulary.
6.2 Lessons for KB construction
The GPTKB prototype reveals several important lessons for LLM-based KB construction. In particular, we find it notable that building such a large KB was possible so quickly, with a relatively small model.
Precision-recall trade-off
The biggest challenge in our view is precision, both in terms of hallucinated entities, and triples. We do not expect that larger models alone will solve this problem, because the long tail, where hallucinations occur, would likely just be pushed farther, but remain difficult to delineate. Tuning the precision-recall tradeoff, for example, via more conservative prompts, or via thresholding based on elicited confidence values Xiong et al. (2023), might be a way forward.
Consolidation challenges
Our exploration surfaced a potpourri of other challenges, some of them known to KB construction research since years Weikum et al. (2021), others requiring novel adaptations in the light of LLMs. These concern, in particular, (1) NER for short labels without context, (2) entity deduplication, (3) entity canonicalization, (4) literal typing and canonicalization, (5) relation clustering, (6) relation organization in terms of subrelations, (7) class clustering, (8) taxonomy construction, and (9) triple verification. While we explored simple techniques for 1, 2, 5, 7 and 8, each has room upward, and the other tasks were not treated so far.
Cost-effectiveness
The GPTKB approach deviates from traditional Wikimedia- and data-integration focused KB construction, and although its precision still needs improvement, it stands out with its potential for cost efficiency. In a back-of-the-envelope calculation, Paulheim (2018) estimated the cost per triple for existing manual KB construction projects at \$2-6, and for existing automated KB construction projects at \$0.01-0.15. In contrast, in our prototypical execution, the cost is just \$0.00009 per correct triple (\$2.2K/24M triples), i.e., over 100x less than with previous methods.
6.3 Other LLMs
We performed a parallel exploration using Llama 3.1 70B on local HPC hardware, but found this not competitive in terms of runtime. We also envisioned a run using the strongest publicly available LLM, GPT-4o, however, by its 15x higher API cost, and its estimated 25x larger parameter set, assuming knowledge is roughly proportional to parameters, at about \$825K, this is beyond our budget.
7 Conclusion
We proposed to build large general-domain KBs entirely from LLMs. We showed that large-scale KB construction from LLMs is feasible, while highlighting specific challenges arising around precision, and KB construction tasks like entity recognition, entity and property canonicalization, and taxonomy construction. As a prototype, we used GPT-4o-mini to construct GPTKB, which contains 105 million assertions for over 2.9 million entities, at a cost 100x cheaper than existing projects.
Our work is a landmark for two fields: For NLP, for the first time, it provides constructive insights into what LLMs know (or believe). For the Semantic Web, it shows novel ways forward for the long stale topic of open-domain KB construction. GPTKB is accessible at http://gptkb.org.
8 Limitations
GPTKB presents a proof of concept that so far does not satisfy the precision requirements of most downstream use cases, nor is its precision even accurately established, because it contains many long-tail triples without web sources that enable confirming or refuting their truth. The knowledge in our KB can intrinsically not be sourced back to any specific document and contributor. Also, as we do not prescribe any standardized vocabulary, and only apply lightweight consolidation, our KB does not yet possess the consistency of existing projects. If you consider using a KB in production, for the time being, consider an alternative like Wikidata.
As there are different ways of how an LLM can be prompted, GPTKB, based at its core on one knowledge elicitation prompt, can only provide one constructive materialization of LLM knowledge, but others with different content and recall or precision exist (see Section 6.1 on “LLM epistemiology”).
Our materialized KB is based on a closed-source LLM that also does not come with guarantees regarding persistent servicing. In the past, similar models have been discontinued, hence, long-term reproducibility is not guaranteed. An execution based on an open model is planned.
Appendix A Further KB Content
Tables 3 and 4 provide further examples of GPTKB content, highlighting strengths as well as weaknesses. For Albuquerque, most triples appear correct, but 2 of the 3 sister cities are not, nor can several of the notablePerson values be verified. In the second case, most triples are correct, but the spouse is made up. In the online OpenReview appendix to this submission, we provide a larger sample of 10K triples.
Relation Object
annualEvent Albuquerque International Balloon Fiesta
area 189.5 square miles
areaCode 505
culturalEvent Albuquerque Film & Music Experience
New Mexico State Fair
Route 66 Summerfest
Albuquerque Pride
Gathering of Nations
elevation 5,312 feet
famousFor New Mexican cuisine
Film industry
Hot air ballooning
founded 1706
governmentType Strong mayor-council
instanceOf City
landmark ABQ BioPark
Old Town Albuquerque
Indian Pueblo Cultural Center
Sandia Peak Tramway
Petroglyph National Monument
locatedIn New Mexico
majorAirport Albuquerque International Sunport
nickname Duke City
notablePerson Larry David
Val Kilmer
Demi Moore
Jessica Chastain
Kurt Russell
Neil Patrick Harris
Billy the Kid
Vince Gilligan
Conan O’Brien
Georgia O’Keeffe
Walter White (fictional character)
Rudy Giuliani
Breaking Bad cast
population 560,000
river Rio Grande
sisterCity Torreón, Mexico
Chihuahua, Mexico
Hiroshima, Japan
sportsTeam Albuquerque Isotopes
New Mexico United
timeZone Mountain Standard Time
university University of New Mexico
zipCode 87101
Table 3: Excerpt of GPTKB triples for subject Albuquerque.
Predicate Object
almaMater Georgia Institute of Technology
birthPlace Panama
children 2
genre Humor
instanceOf Person
knownFor PHD Comics
nationality American
occupation Cartoonist
Filmmaker
Writer
spouse Tanya K. K. Cham
website http://www.phdcomics.com
Table 4: Excerpt of GPTKB triples for subject Jorge Cham of PhDComics.